can you please help me with the math problem i got stuck on.
my teacher says we need the bottom equation to solve the top equation then to find the answer to the bottom equation after.
0.3x-0.2y=0.5
x-2y=-5

Answer by nerdybill(7384)

my teacher says we need the bottom equation to solve the top equation then to find the answer to the bottom equation after.
0.3x-0.2y = 0.5 (top equation)
x- 2y = -5 (bottom equation)
.
Taking the "bottom equation", let's solve for 'x':
x- 2y = -5
adding 2y to both sides we get:
x = 2y -5
.
Now, we take the above, and substitute it into the "top equation":
0.3x-0.2y = 0.5
0.3(2y -5)-0.2y = 0.5
distributing the 0.3 to terms inside the parenthesis:
0.6y-1.5 -0.2y = 0.5
combining y's:
0.4y-1.5 = 0.5
adding 1.5 to both sides:
0.4y = 2
y = 2/0.4
y = 5 (solution for y)
.
Finally, we take the above solution of y and plug it back into the bottom equation and solve for x:
x- 2y = -5
x- 2(5) = -5
x- 10 = -5
adding 10 to both sides:
x = 5 (solution for x)
.
Solution:
y=5
x=5
.
The above method is called the "substitution method".